An object of mass 1kg is falling from the height h=10m.calculate
a) the total energy of an object at h=10m

b) potential energy of the object when it is at h=4m

c) kinetic energy of the object when it is at h=4m

d) what will be the speed of the object when it hits the ground
1
Expert's answer
2021-08-16T08:35:48-0400

Given:

"m=1\:{\rm kg},\quad h=10\: {\rm m},\quad g=9.8\: \rm m/s^2"


a)

"E=E_k+E_p=\frac{mv^2}{2}+mgh""=0+1*9.8*10=98\:\rm J"

b)

"E_p=mgh=1*9.8*4=39\:\rm J"

c)

"E_k=E-E_p=98-39=59\:\rm J"

d)

"E_{kf}=98\:{\rm J}=\frac{mv_f^2}{2}"

"v_f=\sqrt{2*98/1}=14\:\rm m/s"